OpenStudy (anonymous):

what is the solution to this equation? 3x - x + 8 + 5x - 2 = 10 A. x=16/7 B. x=19/9 C. x=4/9 D. x=4/7

OpenStudy (anonymous):

D

OpenStudy (anonymous):

1 Simplify 3x−x+8+5x−2 to 3x−x+5x+6 3x−x+5x+6=10 2 Simplify 3x−x+5x+6 to 7x+6 7x+6=10 3 Subtract 6 from both sides 7x=10−6 4 Simplify 10−6 to 4 7x=4 5 Divide both sides by 7 x=47

OpenStudy (anonymous):

x=4/7 sorry

OpenStudy (anonymous):

thank you here's a medal

OpenStudy (anonymous):

np and thanks